OpenStudy (anonymous):

HELP PLEASE. A person deposited $500 in a saving account that pays 5% annual interest that is compounded yearly, at the end of 10 years , how much money will be in the saving account?

OpenStudy (anonymous):

If you remember the formula: a = p(1 + r/n)^(nt) So we have p = $500, t = 10 years, r = 5% or 0.05, n = 1 (annually compound) So we need to find a (compounded amount) a = 500(1 + 0.05/1)^(1*10) = 500(1.05)^10 = $814.45 which is what you found.